OpenStudy (anonymous):

solve by factoring 8x^2- 10x-25=0 a: x=-0.8 or x=-0.8 c: x=1.25 or x=1.25 b: x=-2.5 or x=-2.5 d: x=-1.25 or x=2.5

OpenStudy (anonymous):

\[8x ^{2}-20x+10x-25=0\] \[4x(2x-5)+5(2x-5)=0\] \[(4x+5)(2x-5)=0\] \[4x+5=0\] \[2x-5=0\] x==-5/4 or 5/2 x=-1.25 or 2.5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you're welcome